The ratio of boys to girls in a class is 5:3. If there are 120 students, how many girls are there? Please show work.
Answer by vicgonzerx(31)

let g=girls
b=boys
the ratio is 5:3=5/3
b/g=5/3
cross mltiply
b=g(5/3)------eq 1
b + g = 120-----eq 2
solve g and b substitution
substitute eq 1-eq 2
g(5/3)+g=120
getting common denominator to add directly
(5g+3g)/3=120
cross multiply
8g=360

then divide it by 8
8g/8=360/8
g=45--------------------girls number
for boys
use eq -1
b=g(5/3)
b=45(5/3)
b=75---------boys number
checking
student number=120
75+45=120-----